OpenStudy (anonymous):

f(x) =sqrt x-5 find f-1

OpenStudy (michele_laino):

please add -1 to both sides of your equation

OpenStudy (kamibug):

Do you mean f(x)-1 or f^-1? :)

OpenStudy (anonymous):

yes f-^1

OpenStudy (michele_laino):

is your function this? \[f(x)=\sqrt{x-5}\]

OpenStudy (anonymous):

i think lol

OpenStudy (kamibug):

Kk, yeah, then you're looking for the inverse. :) Just follow the steps in the link I posted up there. ^.^

OpenStudy (anonymous):

im still confused

OpenStudy (kamibug):

Kay, I'll help. Here are the steps... Step 1: Rewrite f(x) as y. y = sqrt x-5 Step 2: Switch x and y. x = sqrt y-5 Step 3: Solve for y. y = x^2 + 5 Step 4: Rewrite y as f^-1(x). f^-1(x) = x^2 + 5

OpenStudy (anonymous):

A. f-1 (x) = x2 +25; 25 B. f -1 (x) = x2 + 5; y 5 C. f-1 (x) = x2 +25; 5 D. f -1 (x) = (x + 5)2; y 5

OpenStudy (kamibug):

It's B. =]